OpenStudy (anonymous):

Which of the following vectors are orthogonal to (1,5)? Check all that apply a. (-2,5) b. (-5,1) c.(-1,-5) d. 10,-2) I'm thinking b and d, can someone confirm whether or not any of the others work and whether or not I'm right?

OpenStudy (phi):

do you know how to do a dot product? the dot product of 2 vectors that are orthogonal will be zero.

OpenStudy (phi):

for example, <-2,5> dot <1,5>= -2*1 +5*5= -2+25= 23 that is not zero, so not choice A

OpenStudy (anonymous):

Pretty sure it would be "B and D"

OpenStudy (anonymous):

Ah, I think I got it! (-5,1) and (1,5) =-5+5=0 (-1,-5) and (1,5)=-1+-25=-26 (10,-2) and (1,5)=10+-10=0

OpenStudy (jhannybean):

Yeah you got it :) Just make sure you show them as vectors and not points!

OpenStudy (anonymous):

Good job!